Evaluate the limit, if it exists. If not, enter "n" below.
limh→0
(2+h)^3 −8
--------------
h
(the whole thing is divided by h)
1
Expert's answer
2020-09-27T17:38:05-0400
"\lim_{h\to0} \dfrac{(2+h)^3-8}{h} = \\
=\lim_{h\to0} \dfrac{8+12h+6h^2+h^3-8}{h} = \\
=\lim_{h\to0} (12+6h+h^2) = 12"